# How does Watermarking Affect Visual Language Models in Document Understanding?

### v1

@@ 1 Introduction
The rapid advancement of Visual Language Models (VLMs) has transformed document understanding capabilities, enabling sophisticated processing of complex multimodal content without relying on traditional Optical Character Recognition (OCR) technologies. Unlike conventional document processing systems that convert text content into machine-readable characters (Srihari, 1986; Hwang et al., 2021; Hong et al., 2022), modern VLMs directly process original document images, seamlessly integrating visual features with textual information. This OCR-free approach eliminates potential recognition errors in documents with complex layouts, non-standard fonts, and visual noise (Kim et al., 2022; Liu et al., 2024b; Hu et al., 2024b), while preserving critical layout and semantic structures that enhance performance in tasks such as question answering, information retrieval, and cross-page analysis.
However, this advancement introduces new robustness challenges that remain largely unexplored. While considerable research has examined the vulnerability of VLMs to adversarial inputs and visual noise in general contexts, the specific impact of common document modifications, particularly watermarks on VLMs performance represents a critical gap in our understanding. Unlike Large Language Models (LLMs), where robustness mechanisms against harmful queries are relatively well-established, VLMs’ multimodal nature creates unique vulnerability surfaces where image features can potentially bypass or undermine existing safeguards.
@@ 2.1 Visual Language Models for Document Understanding
The field of visual document understanding focuses on understanding structure in visual documents and extracting key information to convert structured or semi-structured forms into machine-readable formats. A visual document is a document that contains a variety of visual and text elements. These documents integrates a variety of visual elements, including paragraphs, charts, tables, etc. Text elements (such as paragraphs and lists) and visual elements (such as charts and tables) together constitute the semantics of the documents. QA is one of the key tasks in visual document understanding, which involves answering natural language questions based on the context information of visual documents(Ding et al., 2024).
With the rapid development of visual document understanding requirements, VLMs technology for visual document understanding is introduced: Qwen-VL (Bai et al., 2023) retain the Q-Former architecture while replacing the original language model with a larger one. In order to extend the capability of VLMs to visual document understanding, mPLUG-DocOwl (Ye et al., 2023a) proposes a modular model based on mPLUG-Owl (Ye et al., 2023b) for document understanding without OCR. These comprehensive approaches highlight the potential of VLMs in handling complex visual language tasks (Li et al., 2024).
@@ 3.1 Document Understanding Task and VLMs Performance Metric
where $\mathbbm{1}(\cdot)$ is the indicator function, which takes the value of 1 if model’s generated answer $f(q,i)$ matches $gt$, and 0 otherwise.
To evaluate the impact of watermark on the model’s performance, we inject a watermark $w$ into each input image $i$ in $\mathcal{D}_{\text{eval}}$, resulting in a perturbed evaluation dataset, denoted as $\mathcal{D}_{\text{eval}}^{\prime}$, which consists of samples in the form of $(q,i+w,gt)$.
@@
The PDR metric directly captures the degradation in model performance when processing watermarked documents compared to original documents, providing a clear measure of how significantly watermarks interfere with the model’s document understanding capabilities. A higher PDR value indicates a stronger disruptive effect of watermark on the model.
@@ 3.2 Experimental Setting
We construct our datasets by randomly sampling 100 images across 3 types of documents, including text, chart, and table from the public dataset of mPLUG-DocOwl2 (Hu et al., 2024a) and some scientific data from National Vital Statistics System website. These images are categorized into four classes: TextS, ChartS, ChartM, and TableS, corresponding to text document images, chart document images, and table document images, along with their respective question types (single choice or multiple response).
@@
This diverse collection of document types and question formats enables us to comprehensively evaluate how watermarks affect VLMs across different document understanding scenarios, from simple text extraction to complex chart interpretation and tabular data analysis.
@@
We select 4 representative sate-of-the-art VLMs for our evaluation: mPLUG-DocOwl2 (Hu et al., 2024b) is designed specifically for document understanding tasks, employing a specialized architecture for processing document layouts and mixed content types.
@@
We primarily investigate the impact of watermark position and content on interference effect. To ensure experimental consistency and isolate the effects of other watermark properties, we keep several key parameters fixed: (1) We set watermark color as black, and set the watermark transparency to 0.5 to ensure that the existence of the watermarks does not affect the visibility of the document to humans, but also has a certain impact on the performance of the model, which is more in line with the realistic scene of document watermarks; (2) We fix watermark angle at 0 degree to eliminate orientation as a variable; (3) Additionally, in each experimental condition, the proportion of watermarks in the total area of the document image is distributed between 10% and 80%.
@@ 4.1.2 Attention Mechanism Analysis
To further explore the underlying reasons why watermarks with different positions produce different interference effects, we calculate the variation in attention weights generated by mPLUG-DocOwl2 model between the watermarked multimodal input sequence and the non-watermarked input sequence, and use it to produce the attention weight variation heatmap (see Figure 3). The visualization reveals distinct patterns: the model adjusts attention weights across the entire image under the Scattered setting (see Figure 3(b)), while attention changes are concentrated in the middle (see Figure 3(a)) and beginning portions (see Figure 3(c)) of the sequence under the Center and Top-left settings, respectively.
@@ 4.2.2 Semantic Interference Analysis
Cross-modal Embedding Cosine Similarity. Cosine similarity evaluates the similarity between two embedding vectors in the semantic space by measuring the directional similarity between them. The higher the cosine similarity score is, the more similar the two embedding vectors are. We compute the cosine similarity between the cross-modal embedding vectors obtained from the watermarked and non-watermarked input sequence using the mPLUG-DocOwl2 model (see Figure 5). The results reveal that the ‘MARK’ watermark leads to the lowest cosine similarity with the original input (0.46), whereas the Mask watermark results in the highest similarity (0.86). This observation further confirms that although the Mask watermark maximizes the occlusion of key information, it induces minimal changes in the cross-modal embeddings’ semantic information. Consequently, this can explain why the average PDR value under the Mask setting is the lowest.
@@ A.1 Impact of Watermark Color
We compare the PDR across datasets with different watermark colors on mPLUG-DocOwl2, all positioned in a Scattered position. The results, shown in Section A.1, indicate that average PDR scores across different colors are similar. This observation can be attributed to how modern VLMs process textual and visual information. These models primarily focus on high-level semantic understanding rather than low-level pixel variations. As a result, changes in watermark color do not significantly affect the model’s interpretation of the document, leading to minimal differences in interference effects.
@@ A.2 Impact of Watermark Area Ratio
Figure 7: PDR before and after JPEG compression.
A.3 Image Preprocessing for Robustness Improvement
@@
A.4 Processing Architecture of LVLMs
@@
Figure 8: The structure of LVLMs for processing input sequences.

### v2

@@ 1 Introduction
The rapid advancement of Visual Language Models (VLMs) has transformed document understanding capabilities, enabling sophisticated processing of complex multimodal content without relying on traditional Optical Character Recognition (OCR) technologies. Unlike conventional document processing systems that convert text content into machine-readable characters (Srihari, 1986; Hwang et al., 2021; Hong et al., 2022), modern VLMs directly process original document images, seamlessly integrating visual features with textual information. This OCR-free approach eliminates potential recognition errors in documents with complex layouts, non-standard fonts, and visual noise (Kim et al., 2022; Liu et al., 2024b; Hu et al., 2024), while preserving critical layout and semantic structures that enhance performance in tasks such as question answering, information retrieval, and cross-page analysis.
However, this advancement introduces new robustness challenges that remain largely unexplored. While considerable research has examined the vulnerability of VLMs to adversarial inputs and visual noise in general contexts, the specific impact of common document modifications, particularly watermarks on VLMs performance, represents a critical gap in our understanding. Unlike Large Language Models (LLMs), where robustness mechanisms against harmful queries are relatively well-established (Mei et al., 2024), VLMs’ multimodal nature creates unique vulnerability surfaces where image features can potentially bypass or undermine existing safeguards.
@@ 2.1 Visual Language Models for Document Understanding
The field of visual document understanding focuses on understanding structure in visual documents and extracting key information to convert structured or semi-structured forms into machine-readable formats. A visual document is a document that contains a variety of visual and text elements. These documents integrate a variety of visual elements, including paragraphs, charts, tables, etc. Text elements (such as paragraphs and lists) and visual elements (such as charts and tables) together constitute the semantics of the documents. QA is one of the key tasks in visual document understanding, which involves answering natural language questions based on the context information of visual documents(Ding et al., 2024).
With the rapid development of visual document understanding requirements, VLMs technology for visual document understanding is introduced: Qwen-VL (Bai et al., 2023) retains the Q-Former architecture while replacing the original language model with a larger one. In order to extend the capability of VLMs to visual document understanding, mPLUG-DocOwl (Ye et al., 2023a) proposes a modular model based on mPLUG-Owl (Ye et al., 2023b) for document understanding without OCR. These comprehensive approaches highlight the potential of VLMs in handling complex visual language tasks (Li et al., 2024).
@@ 3.1 Document Understanding Task and VLMs Performance Metric
where $\mathbbm{1}(\cdot)$ is the indicator function, which takes the value of 1 if the model’s generated answer $f(q,i)$ matches $gt$, and 0 otherwise.
To evaluate the impact of the watermark on the model’s performance, we inject a watermark $w$ into each input image $i$ in $\mathcal{D}_{\text{eval}}$, resulting in a perturbed evaluation dataset, denoted as $\mathcal{D}_{\text{eval}}^{\prime}$, which consists of samples in the form of $(q,i+w,gt)$.
@@
The PDR metric directly captures the degradation in model performance when processing watermarked documents compared to original documents, providing a clear measure of how significantly watermarks interfere with the model’s document understanding capabilities. A higher PDR value indicates a stronger disruptive effect of watermarks on the model.
@@ 3.2 Experimental Setting
We construct our datasets by randomly sampling 100 images for each of the three document types — text, chart, and table — from the public mPLUG-DocOwl2 dataset (Hu et al., 2024) and the LongDocURL dataset available on Hugging Face (Dengchao, 2024). These images are categorized into four classes: TextS, ChartS, ChartM, and TableS, corresponding to text document images, chart document images, and table document images, along with their respective question types (single choice or multiple response).
@@
For text-based documents, 50% are from real estate and finance, and 50% from law-related studies. Chart and table documents are extracted from academic papers in computer science, physics, and related fields. This diverse collection of document types and question formats enables us to comprehensively evaluate how watermarks affect VLMs across different document understanding scenarios, from simple text extraction to complex chart interpretation and tabular data analysis.
@@
We select 4 representative sate-of-the-art VLMs for our evaluation: mPLUG-DocOwl2 (Hu et al., 2024) is designed specifically for document understanding tasks, employing a specialized architecture for processing document layouts and mixed content types.
@@
We primarily investigate the impact of watermark position and content on the interference effect. All watermarks are programmatically added to the document images using the Python Pillow library. To ensure experimental consistency and isolate the effects of other watermark properties, we keep several key parameters fixed: (1) We set watermark color as black and set the watermark transparency to 0.5 to ensure that the existence of the watermarks does not affect the visibility of the document to humans, but also has a certain impact on the performance of the model, which is more in line with the realistic scene of document watermarks; (2) We fix the watermark angle at 0 degrees to eliminate orientation as a variable; (3) Additionally, in each experimental condition, the proportion of watermarks in the total area of the document image is distributed between 10% and 80%. Experiments on watermark angle and opacity are included in the appendix.
@@ 4.1.2 Attention Mechanism Analysis
To further explore the underlying reasons why watermarks with different positions produce different interference effects, we calculate the variation in attention weights generated by mPLUG-DocOwl2 model between the watermarked multimodal input sequence and the non-watermarked input sequence, and use it to produce the attention weight variation heatmap (see Figure 3). The same analysis is also conducted for Qwen2.5-VL and LLaVA-v1.5, with results in the appendix. The visualization reveals distinct patterns: the model adjusts attention weights across the entire image under the Scattered setting (see Figure 3(b)), while attention changes are concentrated in the middle (see Figure 3(a)) and beginning portions (see Figure 3(c)) of the sequence under the Center and Top-left settings, respectively.
@@ 4.2.2 Semantic Interference Analysis
Cross-modal Embedding Cosine Similarity. Cosine similarity evaluates the similarity between two embedding vectors in the semantic space by measuring the directional similarity between them (Zhu et al., 2025; Salton et al., 1975). The higher the cosine similarity score is, the more similar the two embedding vectors are. We compute the cosine similarity between the cross-modal embedding vectors obtained from the watermarked and non-watermarked input sequence using the mPLUG-DocOwl2 model (see Figure 5). The results reveal that the ‘MARK’ watermark leads to the lowest cosine similarity with the original input (0.46), whereas the Mask watermark results in the highest similarity (0.86). This observation further confirms that although the Mask watermark maximizes the occlusion of key information, it induces minimal changes in the cross-modal embeddings’ semantic information. Consequently, this can explain why the average PDR value under the Mask setting is the lowest.
@@ A.1 Impact of Watermark Color
We compare the PDR across datasets with different watermark colors on mPLUG-DocOwl2, all positioned in a Scattered position. The results, shown in Table 3, indicate that average PDR scores across different colors are similar. This observation can be attributed to how modern VLMs process textual and visual information. These models primarily focus on high-level semantic understanding rather than low-level pixel variations. As a result, changes in watermark color do not significantly affect the model’s interpretation of the document, leading to minimal differences in interference effects.
@@ A.2 Impact of Watermark Area Ratio
A.3 Impact of Watermark Opacity
As shown in Table 4, increasing the opacity of the watermark leads to a significant rise in performance degradation, with the average PDR increasing from 8.5% at low opacity ($\alpha=0.2$) to 21.3% at high opacity ($\alpha=0.8$). This trend suggests that higher visual prominence caused by denser watermark opacity results in stronger disruption of model attention. The underlying reason may be that more opaque watermarks dominate the visual space, increasing misalignment in the model’s attention mechanism and impairing its ability to focus on relevant document content.
Table 4: Average PDR(%) for different watermark opacity levels on mPLUG-DocOwl2.
Opacity ($\alpha$) Avg. PDR (%)
0.2 (light) 8.5
0.5 (medium) 15.7
0.8 (dense) 21.3
A.4 Impact of Watermark Rotation
As shown in Table 5, variations in watermark rotation angle have minimal effect on model performance. The average PDR remains relatively stable across different angles, with 14.2% at $0^{\circ}$, 14.5% at $45^{\circ}$, and 14.1% at $90^{\circ}$. This consistency indicates that modern VLMs exhibit strong robustness to moderate geometric transformations in watermark orientation. A possible explanation is that the spatial encoding mechanisms of current vision backbones effectively normalize such rotation variations, making them less disruptive to cross-modal alignment.
Table 5: Average PDR(%) for different watermark rotation angles on mPLUG-DocOwl2.
Rotation Angle Avg. PDR (%)
$0^{\circ}$ 14.2
$45^{\circ}$ (diagonal) 14.5
$90^{\circ}$ (vertical) 14.1
A.5 Attention Heatmap Analysis Across VLMs
As shown in the Figure 7 and Figure 8, these are the attention variation heatmaps of Qwen2.5-VL and LLaVA-v1.5. It is evident that the Qwen2.5-VL exhibits stronger robustness compared to LLaVA-v1.5 and mPLUG-DocOwl2, which is consistent with the experimental results.
Figure 7: Heatmap of attention wight variation between watermarked multimodal input sequences and original one for Qwen2.5-VL.
Figure 8: Heatmap of attention wight variation between watermarked multimodal input sequences and original one for LLaVA-v1.5.
A.6 Image Preprocessing for Robustness Improvement
@@
Figure 9: PDR before and after JPEG compression.
A.7 Processing Architecture of VLMs
@@
Figure 10: The structure of VLMs for processing input sequences.